# The radius of statistical efficiency

### v1

@@ 1.2 Notation
Throughout, we let $\mathbf{R}^{d}$ denote the standard $d$-dimensional Euclidean space with dot product $\langle x,y\rangle=x^{\top}y$ and norm $\|x\|_{2}=\sqrt{\langle x,x\rangle}$. The unit sphere in $\mathbf{R}^{d}$ will be denoted by $\mathbb{S}^{d-1}$, while the set of nonnegative vectors will be written as $\mathbf{R}^{d}_{+}$. The symbol $\mathbf{R}^{m\times n}$ will denote the Euclidean space of real $m\times n$ matrices, endowed with the trace inner product $\langle X,Y\rangle=\operatorname{tr}(X^{\top}Y)$. The symbol $\otimes$ denotes Kronecker product. The Frobenius and operator norms will be written as $\|\cdot\|_{F}$ and $\|\cdot\|_{\rm op}$, respectively. The singular values of a matrix $A\in\mathbf{R}^{m\times n}$ will be arranged in nonincreasing order:
$\sigma_{1}(A)\geq\sigma_{2}(A)\geq\ldots\geq\sigma_{m\wedge n}(A).$
The space of real symmetric $d\times d$ matrices is denoted by $\bm{S}^{d}$ and is equipped with the trace product as well. The cone of $d\times d$ positive semidefinite matrices will be written as $\bm{S}^{d}_{+}$. The eigenvalues of a matrix $A\in\bm{S}^{d}$ will be arranged in nonincreasing order:
$\lambda_{1}(A)\geq\lambda_{2}(A)\geq\ldots\geq\lambda_{d}(A).$
For any subspace $\mathcal{K}\subset\mathbf{R}^{d}$, the symbol $\mathsf{P}_{\mathcal{K}}\colon{\mathbf{R}}^{d}\rightarrow\mathcal{K}$ will denote the orthogonal projection onto $\mathcal{K}$. The compression of any matrix $A\in{\mathbf{R}}^{d\times d}$ to $\mathcal{K}$, denoted $A|_{\mathcal{K}}\colon\mathcal{K}\rightarrow\mathcal{K}$, is the map $A|_{\mathcal{K}}:=\mathsf{P}_{\mathcal{K}}A\mathsf{P}_{\mathcal{K}}$.
@@
We will require some background on the Wasserstein geometry on the space of probability measures on $\mathbf{R}^{d}$. In order to streamline the reading, we record here only the most essential notation that we will need. A detailed review of Wasserstein geometry appears in Section A. To this end, we let $\mathcal{P}_{p}(\mathbf{R}^{d})$ be the space of measures $\mu$ on $\mathbf{R}^{d}$ with a finite $p$’th moment $\operatorname{\mathbb{E}}_{\mu}\|x\|^{p}_{p}<\infty$. The subset of measures of $\mathcal{P}_{p}(\mathbf{R}^{d})$ that are centered, meaning $\operatorname{\mathbb{E}}_{\mu}[x]=0$, will be written as $\mathcal{P}_{p}^{\circ}(\mathbf{R}^{d})$. When the space $\mathbf{R}^{d}$ is clear from context, we use the shorthand $\mathcal{P}_{p}$ and $\mathcal{P}_{p}^{\circ}$. A convenient metric on $\mathcal{P}_{p}$ is furnished by the Wasserstein distance $W_{p}(\mu,\nu)$; see Section A for details. The distance function to a set of measures $Q\subset\mathcal{P}_{p}$ is defined by $W_{p}(\mu,Q)=\inf_{\nu\in Q}W_{p}(\mu,\nu)$. The symbol $\Sigma_{\mu}=\operatorname{\mathbb{E}}_{\mu}xx^{\top}$ will denote the second moment matrix of any measure $\mu\in\mathcal{P}_{2}$. In the rest of the paper, we will use the symbol $\mathsf{D}$ to denote a distinguished measure associated with the problem of interest, while we use $\mu$ as a placeholder for arbitrary measures.
@@ 2 The distance to ill-conditioned problems
In this section, we formally define the radius of statistical efficiency (RSE), and develop some techniques for computing it. Throughout, we will focus on the stochastic optimization problem
@@
$\min_{\beta\in\mathcal{M}}~{}f(\beta)\quad\text{where}\quad f(\beta)=\mathop{\operatorname{\mathbb{E}}}_{z\sim\mathsf{D}}\ell(\beta;z).$ $\rm{SP}(\mathsf{D})$
@@
Here, the set $\mathcal{M}\subset{\mathbf{R}}^{d}$ is a $C^{2}$ manifold and $z$ is drawn from a distribution $\mathsf{D}\in\mathcal{P}_{2}(\mathcal{Z})$, where $\mathcal{Z}$ is a finite-dimensional Euclidean space equipped with its Borel $\sigma$-algebra. We assume that the function $\ell(\beta;z)$ is measurable and twice differentiable in $\beta$ for every $z$ and that $f$ is $C^{2}$-smooth. We also make the blanket assumption that the gradient and Hessian of $\ell(\cdot;z)$ are $\mathsf{D}$-integrable.
@@
The difficulty of solving the problem $\rm{SP}(\mathsf{D})$ from finitely many samples $z_{1},z_{2},\ldots,z_{n}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\mathsf{D}$ is tightly controlled by a matrix akin to Fisher Information. This object, which we now describe, plays a central role in our work. Let ${\beta}_{\star}$ be a minimizer of $\rm{SP}(\mathsf{D})$ and define the solution map:
@@
$\sigma(v)=\operatornamewithlimits{argmin}_{\beta\in\mathcal{M}\cap B_{\varepsilon}({\beta}_{\star})}~{}f(\beta)-\langle v,\beta\rangle.$
@@
Thus, the set $\sigma(v)$ is comprised of all solutions to a problem obtained from $\rm{SP}(\mathsf{D})$ by subtracting a linear/tilt perturbation $\langle v,\beta\rangle$. Clearly, a desirable property is for $\sigma$ to be single-valued and smooth. With this in mind, we introduce the following notion due to Poliquin and Rockafellar [50].
@@
Definition: (Tilt-stable minimizer) The point ${\beta}_{\star}$ is a tilt-stable minimizer of $\rm{SP}(\mathsf{D})$ if the map $\sigma(\cdot)$ satisfies $\sigma(0)={\beta}_{\star}$ and is single-valued and $C^{1}$-smooth on some neighborhood of the origin. Then the regularity modulus of the problem is defined to be ${\rm REG}(\mathsf{D})=\|\nabla\sigma(0)\|_{\rm op}.$ If ${\beta}_{\star}$ is not tilt-stable, we call ${\beta}_{\star}$ unstable and set ${\rm REG}(\mathsf{D})=+\infty$.
@@
In particular, we will regard ${\rm REG}(\mathsf{D})$ as the measure of difficulty of solving the problem $\rm{SP}(\mathsf{D})$. When $\mathcal{M}$ is the whole space, ${\beta}_{\star}$ is a tilt-stable minimizer if and only if the Hessian $\nabla^{2}f({\beta}_{\star})$ is nonsingular, in which case equality $\nabla\sigma(0)=[\nabla^{2}f({\beta}_{\star})]^{-1}$ holds [50, Proposition 1.2]. In particular, when $\rm{SP}(\mathsf{D})$ corresponds to maximum likelihood estimation, the matrix $\nabla\sigma(0)$ reduces to the inverse of the Fisher information. More generally, tilt-stability can be characterized either in terms of definiteness of the covariant Hessian or the Hessian of the Lagrangian on the tangent space to $\mathcal{M}$. Since we will use both of these viewpoints, we review them now. The reader may safely skip this discussion during the first reading since it will not be used until the appendix.
@@
Let $G=0$ be the local defining equations for $\mathcal{M}$ around ${\beta}_{\star}$. That is $G\colon\mathbf{R}^{d}\to\mathbf{R}^{m}$ is a $C^{2}$-smooth map with surjective Jacobian $\nabla G({\beta}_{\star})$ and such that the two sets $\mathcal{M}$ and $\{\beta:G(\beta)=0\}$ coincide near ${\beta}_{\star}$. Then the tangent space to $\mathcal{M}$ at ${\beta}_{\star}$ is $\mathcal{T}={\rm Null}(\nabla G({\beta}_{\star}))$. Define the Lagrangian function
@@
$\mathcal{I}({\beta}_{\star},\mathsf{D}):=P_{\mathcal{T}}\cdot\nabla^{2}_{\beta\beta}\mathcal{L}({\beta}_{\star},\lambda_{\star})\cdot P_{\mathcal{T}},$ (7)
where $P_{\mathcal{T}}$ is the orthogonal projection onto the tangent space $\mathcal{T}$. If ${\beta}_{\star}$ is a local minimizer of the problem, then $\mathcal{I}({\beta}_{\star},\mathsf{D})$ is positive semidefinite. Conversely:
@@
Moreover, in this case equality $\nabla\sigma(0)=\mathcal{I}({\beta}_{\star},\mathsf{D})^{\dagger}$ holds, where $\dagger$ denotes the Moore–Penrose inverse. In particular, one may regard $\nabla\sigma(0)$ as akin to the inverse of the Fisher information matrix for MLE. Note that when ${\beta}_{\star}$ is tilt-stable, the reciprocal of the regularity modulus $1/{\rm REG}(\mathsf{D})$ coincides the minimal nonzero eigenvalue of $\mathcal{I}({\beta}_{\star},\mathsf{D})$.
@@
Often, the defining equations of the manifold are either unknown or difficult to work with. In this case, tilt-stability can be characterized through second-order expansions along curves. Namely, for any tangent vector $u\in{\mathcal{T}}$, there exists a $C^{2}$-smooth curve $\gamma_{u}\colon(-\epsilon,\epsilon)\to\mathcal{M}$ for some $\epsilon>0$ satisfying $\gamma_{u}(0)={\beta}_{\star}$ and $\dot{\gamma_{u}}(0)=u$. The covariant Hessian of $f$ at ${\beta}_{\star}$ is the unique symmetric bilinear form $\nabla^{2}_{\mathcal{M}}f({\beta}_{\star})\colon{\mathcal{T}}\times{\mathcal{T}}\to\mathbf{R}$ satisfying
$\nabla^{2}_{\mathcal{M}}f({\beta}_{\star})[u,u]=(f\circ\gamma_{u})^{\prime\prime}(0).$
It is classically known that equality holds
@@
where the matrix $\mathcal{I}({\beta}_{\star},\mathsf{D})$ is defined in (7). Consequently, $\nabla^{2}_{\mathcal{M}}f({\beta}_{\star})$ is positive semidefinite when ${\beta}_{\star}$ is a local minimizer. Conversely:
$\nabla^{2}_{\mathcal{M}}f({\beta}_{\star})$ is positive definite on $\mathcal{T}$ if and only if ${\beta}_{\star}$ is a tilt-stable minimizer.
In this case, identifying $\nabla^{2}_{\mathcal{M}}f({\beta}_{\star})$ with a matrix, equality $\nabla\sigma(0)=(P_{\mathcal{T}}\nabla^{2}_{\mathcal{M}}f({\beta}_{\star})P_{\mathcal{T}})^{\dagger}$ holds. In particular, the equality ${\rm REG}(\mathsf{D})^{-1}=\lambda_{\min}(\nabla^{2}_{\mathcal{M}}f({\beta}_{\star}))$ holds.
The sensitivity matrix $\nabla\sigma(0)$ figures prominently in the asymptotic performance of estimation procedures. Notably, building on classical ideas due to Hájek and Le Cam, the recent paper of Duchi and Ruan [35] established a lower bound on asymptotic covariance of arbitrary estimators $\widehat{\beta}_{n}$ of ${\beta}_{\star}$ using $n$ samples $z_{1}\dots,z_{n}.$ The precise lower bound is quite technical, and we refer the interested reader to their paper. In summary, their result shows that if ${\beta}_{\star}$ is a tilt-stable minimizer, then the asymptotic covariance of $\sqrt{n}(\widehat{\beta}_{n}-{\beta}_{\star})$ is lower bounded in the Loewner order by the matrix
$\mathsf{\Sigma}:=\nabla\sigma(0)\cdot\operatorname{Cov}\left(\nabla\ell({\beta}_{\star},z)\right)\cdot\nabla\sigma(0),$ (8)
Moreover, in typical settings the expression in (8) simplifies to the equality $\mathsf{\Sigma}=\nabla\sigma(0)$; this is the case for example for (quasi) maximum likelihood estimation (Section 4) and rank-one matrix regression problems (Section 5). Thus, asymptotically the best error that any estimator can achieve in the direction $u$ is on the order $n^{-1/2}\cdot u^{\top}\mathsf{\Sigma}u$. The direction $u$ with the worst error matches the top eigenvector of $\mathsf{\Sigma}$ and the number of samples necessary to find an accurate approximation of ${\beta}_{\star}$ grows with $\lambda_{\max}(\mathsf{\Sigma})$. Reassuringly, typical algorithms such as sample average approximation [65] and the stochastic projected gradient method [51, 26] match the lower-bound (8) and thus are asymptotically optimal.
@@
Definition: Radius of statistical efficiency (unsupervised) Consider the problem $\mathtt{SP}(\mathsf{D})$ and let $\mathcal{Q}\subset\mathcal{P}_{2}(\mathcal{Z})$ be a distinguished set of distributions. Define the set of ill-conditioned distributions as =$\mathcal{E}=\left\{\mathsf{D}^{\prime}\in\mathcal{Q}\mid\text{There exists a minimizer of ${\rm SP}(\mathsf{D}^{\prime})$ that is unstable}\right\}.$ (9) The radius of statistical efficiency of $\mathsf{D}$ is defined to be $\operatorname{RSE}(\mathsf{D}):=W_{2}(\mathsf{D},\mathcal{E}).$
@@
Problems of supervised learning are distinctly different. The data consists of pairs $z=(x,y)\sim\mathsf{D}$, where $x\sim\mathsf{D}_{x}$ are called feature vectors and $y\sim\mathsf{D}_{y\mid x}$ are the labels. We will assume that the conditional distribution $\mathsf{D}_{y\mid x}$ depends on the features $x$ and a latent parameter $\beta^{\star}$. A typical example is the setting of regression under a model $y=g(x,{\beta}_{\star})+\varepsilon$ where $\varepsilon$ is a noise vector that is independent of $x$. The goal of the corresponding optimization problem $\rm{SP}(\mathsf{D})$ is to recover $\beta^{\star}$. In contrast to unsupervised learning, the latent parameter ${\beta}_{\star}$ is fixed a priori and is not a function of the data distribution. Therefore a natural measure of robustness of the problem is the size of the smallest perturbation to the feature vectors $W_{2}(\mathsf{D}^{\prime}_{x},\mathsf{D}_{x})$ so that the problem ${\rm SP}(\mathsf{D}^{\prime}_{x}\times{\mathsf{D}}_{y\mid x})$ has an unstable minimizer. Notice that the conditional distributions of $y$ given $x$ coincide for the two measures $\mathsf{D}$ and $\mathsf{D}^{\prime}=\mathsf{D}^{\prime}_{x}\times\mathsf{D}_{y\mid x}$. This type of shift exclusively only in the feature data appears often in the literature under the name of covariate shift [61, 52, 1, 69, 54].
@@
Definition: Radius of statistical efficiency (supervised) Consider the problem $\mathtt{SP}(\mathsf{D})$ with $\mathsf{D}=\mathsf{D}_{x}\times\mathsf{D}_{y\mid x}$ and let ${\beta}_{\star}$ be its minimizer. Let $\mathcal{Q}\subset\mathcal{P}_{2}$ be a distinguished set of distributions. Define the set of ill-conditioned distributions as =$\mathcal{E}=\left\{\mathsf{D}^{\prime}_{x}\in\mathcal{Q}\mid\text{${\beta}_{\star}$ is not a tilt-stable minimizer of ${\rm SP}(\mathsf{D}^{\prime}_{x}\times\mathsf{D}_{y\mid x})$}\right\}.$ (10) The radius of statistical efficiency of $\mathsf{D}$ is defined to be $\operatorname{RSE}(\mathsf{D}):=W_{2}(\mathsf{D}_{x},\mathcal{E}).$
@@
Evidently, the quantity ${\rm RSE}(\mathsf{D})$ measures the robustness of the problem because it quantifies the size of a neighborhood around $\mathsf{D}$ for which all problem instances are stable. There is a small nuance in formalizing this statement due to a lack of compactness in the Wasserstein space. Namely, we have to impose the minor assumption that any sequence of measures $\nu_{i}$ for which the problem ${\rm SP}(\nu_{i})$ becomes progressively harder (${\rm REG}(\nu_{i})\to\infty$) must approach the set of ill-conditioned distributions (${\rm RSE}(\nu_{i})\to 0$). In all examples we consider, this holds at least on bounded sets $\mathcal{Q}^{\prime}\subset\mathcal{P}_{2}$. The proof of this elementary observation appears in Appendix B.1.

### v2

@@ 1.2 Notation
Throughout, we let $\mathbf{R}^{d}$ denote the standard $d$-dimensional Euclidean space equipped with the dot product $\langle x,y\rangle=x^{\top}y$ and the induced norm $\|x\|_{2}=\sqrt{\langle x,x\rangle}$ (the $\ell_{2}$ norm). The unit sphere in $\mathbf{R}^{d}$ will be denoted by $\mathbb{S}^{d-1}$, while the set of nonnegative vectors will be written as $\mathbf{R}^{d}_{+}$. The symbol $\mathbf{R}^{m\times n}$ will denote the Euclidean space of real $m\times n$ matrices, equipped with the trace inner product $\langle X,Y\rangle=\operatorname{tr}(X^{\top}Y)$ and the induced norm $\|X\|_{F}=\sqrt{\langle X,X\rangle}$ (the Frobenius norm). The operator norm will be written as $\|\cdot\|_{\rm op}$. The symbol $\otimes$ denotes the Kronecker product. The singular values of a matrix $A\in\mathbf{R}^{m\times n}$ will be arranged in nonincreasing order:
$\sigma_{1}(A)\geq\sigma_{2}(A)\geq\cdots\geq\sigma_{m\wedge n}(A).$
The space of real symmetric $d\times d$ matrices will be denoted by $\bm{S}^{d}$ and is equipped with the trace product as well. The cone of real positive semidefinite (PSD) $d\times d$ matrices will be written as $\bm{S}^{d}_{+}$. The eigenvalues of a matrix $A\in\bm{S}^{d}$ will be arranged in nonincreasing order:
$\lambda_{1}(A)\geq\lambda_{2}(A)\geq\cdots\geq\lambda_{d}(A).$
For any subspace $\mathcal{K}\subset\mathbf{R}^{d}$, the symbol $\mathsf{P}_{\mathcal{K}}\colon{\mathbf{R}}^{d}\rightarrow\mathcal{K}$ will denote the orthogonal projection onto $\mathcal{K}$. The compression of a linear operator $A$ on $\mathbf{R}^{d}$ to $\mathcal{K}$, denoted $A|_{\mathcal{K}}\colon\mathcal{K}\rightarrow\mathcal{K}$, is the restriction of $\mathsf{P}_{\mathcal{K}}A$ to $\mathcal{K}$.
@@
We will require some background on the Wasserstein geometry on the space of probability measures on a Euclidean space $(\mathbf{E},\|\cdot\|)$ equipped with its Borel $\sigma$-algebra. In order to streamline the reading, we record here only the most essential notation that we will need. A more detailed review of Wasserstein geometry appears in Section A. To this end, we let $\mathcal{P}_{p}(\mathbf{E})$ be the space of measures $\mu$ on $\mathbf{E}$ with finite $p$’th moment ${\operatorname{\mathbb{E}}}_{\mu}\|x\|^{p}<\infty$. The subset comprised of measures $\mu\in\mathcal{P}_{p}(\mathbf{E})$ that are centered, meaning $\operatorname{\mathbb{E}}_{\mu}[x]=0$, will be written as $\mathcal{P}_{p}^{\circ}(\mathbf{E})$. When the space $\mathbf{E}$ is clear from context, we use the shorthand $\mathcal{P}_{p}$ and $\mathcal{P}_{p}^{\circ}$. A convenient metric on $\mathcal{P}_{p}$ is furnished by the Wasserstein-$p$ distance
$W_{p}(\mu,\nu)=\min_{\pi\in\Pi(\mu,\nu)}\bigg(\mathop{\operatorname{\mathbb{E}}}_{(x,y)\sim\pi}\|x-y\|^{p}\bigg)^{1/p}$
(see Section A for details). The distance function to a set of measures $\mathcal{Q}\subset\mathcal{P}_{p}$ is defined by $W_{p}(\mu,\mathcal{Q})=\inf_{\nu\in\mathcal{Q}}W_{p}(\mu,\nu)$. The symbol $\Sigma_{\mu}=\operatorname{\mathbb{E}}_{\mu}[xx^{\top}]$ will denote the second-moment matrix of any measure $\mu\in\mathcal{P}_{2}(\mathbf{R}^{d})$. In the rest of the paper, we will use the symbol $\mathsf{D}$ to denote a distinguished measure associated with the problem of interest, while we use $\mu$ as a placeholder for arbitrary measures.
@@ 2 The distance to ill-conditioned problems
In this section, we formally define the radius of statistical efficiency (RSE) and develop some techniques for computing it. Throughout, we will focus on the stochastic optimization problem
@@
$\min_{\beta\in\mathcal{M}}\,f(\beta)\quad\text{where}\quad f(\beta)=\mathop{\operatorname{\mathbb{E}}}_{z\sim\mathsf{D}}\ell(\beta;z).$ $\rm{SP}(\mathsf{D})$
@@
Here the set $\mathcal{M}$ is a $C^{2}$ embedded submanifold of $\mathbf{R}^{d}$ and $z$ is drawn from a distribution $\mathsf{D}\in\mathcal{P}_{2}(\mathcal{Z})$, where $\mathcal{Z}$ is a Euclidean space equipped with its Borel $\sigma$-algebra. We assume that the function $f$ is $C^{2}$-smooth on $\mathcal{M}$, i.e., at each point $\beta\in\mathcal{M}$, there exists a $C^{2}$-smooth local extension of $f$ (which we shall also denote by $f$) defined on an open neighborhood of $\beta$ in $\mathbf{R}^{d}$.
@@
The difficulty of solving the problem $\rm{SP}(\mathsf{D})$ from finitely many samples $z_{1},\ldots,z_{n}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\mathsf{D}$ is tightly controlled by a matrix akin to Fisher information. This object, which we now describe, plays a central role in our work. Let $\beta$$_{\star}$ be a minimizer of $\rm{SP}(\mathsf{D})$ and define the solution map
@@
$\sigma(v)=\operatornamewithlimits{argmin}_{\beta\in\mathcal{M}\cap B_{\varepsilon}({\beta}_{\star})}\,f(\beta)-\langle v,\beta\rangle$
@@
for some $\varepsilon>0$. Thus, ${\beta}_{\star}\in\sigma(0)$ and the set $\sigma(v)$ is comprised of all local minimizers ($\varepsilon$-close to $\beta$$_{\star}$) of the problem obtained from $\rm{SP}(\mathsf{D})$ by subtracting a linear/tilt perturbation $\langle v,\beta\rangle$. Clearly, a desirable property is for $\sigma$ to be single-valued and smooth. With this in mind, we introduce the following notion due to Poliquin and Rockafellar [53].
@@
Definition: (Tilt-stable minimizer) The point $\beta$$_{\star}$ is a tilt-stable minimizer of $\rm{SP}(\mathsf{D})$ if for some $\varepsilon>0$, the map $\sigma(\cdot)$ satisfies $\sigma(0)={\beta}_{\star}$ and is single-valued and $C^{1}$-smooth on some neighborhood of the origin. Then the regularity modulus of the problem is defined to be ${\rm REG}(\mathsf{D})=\|\nabla\sigma(0)\|_{\rm op}.$ If $\beta$$_{\star}$ is not tilt-stable, we call $\beta$$_{\star}$ unstable and set ${\rm REG}(\mathsf{D})=+\infty$.
@@
In particular, we will regard ${\rm REG}(\mathsf{D})$ as a measure of difficulty of solving the problem $\rm{SP}(\mathsf{D})$. If $\mathcal{M}$ is the whole space $\mathbf{R}^{d}$, then $\beta$$_{\star}$ is a tilt-stable minimizer if and only if the Hessian $\nabla^{2}f({\beta}_{\star})$ is nonsingular, in which case equality $\nabla\sigma(0)=[\nabla^{2}f({\beta}_{\star})]^{-1}$ holds [53, Proposition 1.2]. In particular, when $\rm{SP}(\mathsf{D})$ corresponds to maximum likelihood estimation, the matrix $\nabla\sigma(0)$ reduces to the inverse of the Fisher information. More generally, tilt-stability can be characterized either in terms of definiteness of the covariant Hessian or the Hessian of the Lagrangian on the tangent space of $\mathcal{M}$. Since we will use both of these viewpoints, we review them now; for further details regarding the following Lagrangian and intrinsic characterizations, we refer the reader to [42, 53, 26]. The reader may safely skip this discussion during the first reading since it will not be used until the appendix.
@@
Let $G=0$ be the local defining equations for $\mathcal{M}$ around $\beta$$_{\star}$. That is, $G\colon\mathbf{R}^{d}\to\mathbf{R}^{m}$ is a $C^{2}$-smooth map with surjective Jacobian $\nabla G({\beta}_{\star})$ and the two sets $\mathcal{M}$ and $\{\beta\in\mathbf{R}^{d}\mid G(\beta)=0\}$ coincide near $\beta$$_{\star}$. Then $\mathcal{T}:=\ker(\nabla G({\beta}_{\star}))$ is the tangent space of $\mathcal{M}$ at $\beta$$_{\star}$. Define the Lagrangian function
@@
$\mathcal{I}({\beta}_{\star},\mathsf{D}):=\mathsf{P}_{\mathcal{T}}\nabla^{2}_{\beta\beta}\mathcal{L}({\beta}_{\star},\lambda_{\star})\mathsf{P}_{\mathcal{T}},$ (7)
where we let $\mathsf{P}_{\mathcal{T}}$ denote the $d\times d$ matrix representation of the orthogonal projection of $\mathbf{R}^{d}$ onto the tangent space $\mathcal{T}$ relative to the standard basis. Since $\beta$$_{\star}$ is a local minimizer of $\rm{SP}(\mathsf{D})$, it follows that $\mathcal{I}({\beta}_{\star},\mathsf{D})$ is positive semidefinite. Conversely:
@@
An application of the implicit function theorem, in tandem with classical results on tilt-stability [42], yields $\nabla\sigma(0)=\mathcal{I}({\beta}_{\star},\mathsf{D})^{\dagger}$ where $\dagger$ denotes the Moore–Penrose pseudoinverse; see, for instance, [26, Theorem 2.7]. In particular, one may regard $\nabla\sigma(0)$ as akin to the inverse of the Fisher information matrix for MLE. Note that when $\beta$$_{\star}$ is tilt-stable, the reciprocal of the regularity modulus ${\rm REG}(\mathsf{D})^{-1}$ coincides with the minimal nonzero eigenvalue of $\mathcal{I}({\beta}_{\star},\mathsf{D})$.
@@
The local defining equations for the manifold $\mathcal{M}$ are often either unknown or difficult to work with. In this case, tilt-stability can be characterized through second-order expansions along curves. Namely, for any tangent vector $u\in{\mathcal{T}}$, there exists a $C^{2}$-smooth curve $\gamma_{u}\colon(-\epsilon,\epsilon)\to\mathcal{M}$ for some $\epsilon>0$ satisfying $\gamma_{u}(0)={\beta}_{\star}$ and $\dot{\gamma_{u}}(0)=u$. Since $\beta_{\star}$ is a critical point of $f$ on $\mathcal{M}$, the covariant Hessian of $f$ at $\beta$$_{\star}$ is the unique symmetric bilinear form $\nabla^{2}_{\mathcal{M}}f({\beta}_{\star})\colon{\mathcal{T}}\times{\mathcal{T}}\to\mathbf{R}$ satisfying
$\nabla^{2}_{\mathcal{M}}f({\beta}_{\star})[u,u]=(f\circ\gamma_{u})^{\prime\prime}(0)\qquad\forall u\in\mathcal{T}.$
It is classically known that the following equality holds:
@@
where the matrix $\mathcal{I}({\beta}_{\star},\mathsf{D})$ is defined in (7). Thus, viewing $\nabla^{2}_{\mathcal{M}}f({\beta}_{\star})$ as a linear operator on $\mathcal{T}$} (using the Riemannian metric induced by the inclusion $\mathcal{M}\hookrightarrow\mathbf{R}^{d}$), we may identify $\nabla^{2}_{\mathcal{M}}f({\beta}_{\star})$ with the compression of $\mathcal{I}({\beta}_{\star},\mathsf{D})$ to $\mathcal{T}$}. Consequently, $\nabla^{2}_{\mathcal{M}}f({\beta}_{\star})$ is positive semidefinite since $\beta$$_{\star}$ is a local minimizer of $\rm{SP}(\mathsf{D})$. Conversely:
$\nabla^{2}_{\mathcal{M}}f({\beta}_{\star})$ is positive definite if and only if $\beta$$_{\star}$ is a tilt-stable minimizer.
In this case, the equality ${\rm REG}(\mathsf{D})^{-1}=\lambda_{\min}(\nabla^{2}_{\mathcal{M}}f({\beta}_{\star}))$ holds.
The sensitivity matrix $\nabla\sigma(0)$ figures prominently in the asymptotic performance of estimation procedures. Notably, building on classical ideas due to Hájek and Le Cam, the recent paper of Duchi and Ruan [35] established a lower bound on asymptotic covariance of arbitrary estimators $\widehat{\beta}_{n}$ of $\beta$$_{\star}$ using $n$ samples $z_{1},\ldots,z_{n}.$ The precise lower bound is quite technical, and we refer the interested reader to their paper. In summary, their result shows that if $\beta$$_{\star}$ is a tilt-stable minimizer, then the asymptotic covariance of $\sqrt{n}(\widehat{\beta}_{n}-{\beta}_{\star})$ is lower-bounded in the Loewner order by the matrix
$\mathsf{\Sigma}:=\nabla\sigma(0)\cdot\operatorname{Cov}(\nabla\ell({\beta}_{\star};z))\cdot\nabla\sigma(0).$ (8)
Moreover, for the problems considered in this work the expression in (8) simplifies to $\mathsf{\Sigma}\asymp\nabla\sigma(0)$ (equality up to a scalar factor); this is the case for example for (quasi) maximum likelihood estimation (Section 4) and rank-one matrix regression problems (Section 5). Thus, asymptotically the best error that any estimator can achieve in the direction $u$ is on the order $n^{-1/2}\cdot u^{\top}\mathsf{\Sigma}u$. The direction $u$ with the worst error matches the top eigenvector of $\mathsf{\Sigma}$ and the number of samples necessary to find an accurate approximation of $\beta$$_{\star}$ grows with $\lambda_{\max}(\mathsf{\Sigma})$. Reassuringly, typical algorithms such as sample average approximation [69] and the stochastic projected gradient method [54, 26] match the lower bound (8) and thus are asymptotically optimal.
@@
[…]
@@
Problems of supervised learning are distinctly different. The data consists of pairs $z=(x,y)\sim\mathsf{D}$ comprised of feature data $x\sim\mathsf{D}_{x}$ and labels $y\sim\mathsf{D}_{y\mid x}$. We will assume that the conditional distribution $\mathsf{D}_{y\mid x}$ depends on the features $x$ and a latent parameter $\beta_{\star}$. A typical example is the setting of regression under a model $y=g(x,{\beta}_{\star})+\varepsilon$ where $\varepsilon$ is a random noise variable that is independent of $x$. The goal of the corresponding optimization problem $\rm{SP}(\mathsf{D})$ is to recover $\beta$$_{\star}$. In contrast to unsupervised learning, the latent parameter $\beta$$_{\star}$ is fixed a priori and is not a function of the data distribution. Therefore a natural measure of robustness of the problem is the size of the smallest perturbation to the feature vectors $W_{2}(\mathsf{D}^{\prime}_{x},\mathsf{D}_{x})$ so that the problem ${\rm SP}(\mathsf{D}^{\prime}_{x}\times{\mathsf{D}}_{y\mid x})$ has an unstable minimizer. Notice that the conditional distributions of $y$ given $x$ coincide for the two measures $\mathsf{D}$ and $\mathsf{D}^{\prime}=\mathsf{D}^{\prime}_{x}\times\mathsf{D}_{y\mid x}$. This type of shift exclusively only in the feature data appears often in the literature under the name of covariate shift [64, 55, 1, 73, 57].
@@
Definition: Radius of statistical efficiency (supervised) Consider the problem $\mathtt{SP}(\mathsf{D})$ with $\mathsf{D}=\mathsf{D}_{x}\times\mathsf{D}_{y\mid x}$ and let $\beta$$_{\star}$ be its minimizer. Let $\mathcal{Q}\subset\mathcal{P}_{2}$ be a distinguished set of permissible distributions. Define the set of ill-conditioned distributions as $\mathcal{E}:=\left\{\mathsf{D}^{\prime}_{x}\in\mathcal{Q}\mid\text{${\beta}_{\star}$ is not a tilt-stable minimizer of ${\rm SP}(\mathsf{D}^{\prime}_{x}\times\mathsf{D}_{y\mid x})$}\right\}\!.$ (10) The radius of statistical efficiency of $\mathsf{D}$ is defined to be $\operatorname{RSE}(\mathsf{D}):=W_{2}(\mathsf{D}_{x},\mathcal{E}).$
@@
In the supervised setting, we will sometimes write $\operatorname{RSE}(\mathsf{D}_{x})$ in place of $\operatorname{RSE}(\mathsf{D})$, and, similarly, ${\rm SP}(\mathsf{D}_{x})$ and ${\rm REG}(\mathsf{D}_{x})$ in place of ${\rm SP}(\mathsf{D})$ and ${\rm REG}(\mathsf{D})$, respectively. In either setting, the quantity ${\rm RSE}(\mathsf{D})$ evidently measures the robustness of the problem because it quantifies the size of a neighborhood around $\mathsf{D}$ for which all problem instances are stable. There is a small nuance in formalizing this statement due to a lack of compactness in the Wasserstein space; specifically, closed balls in $\mathcal{P}_{2}$ are not compact. To navigate this difficulty, we impose the minor assumption that any sequence of measures $\nu_{i}$ for which the problem ${\rm SP}(\nu_{i})$ becomes progressively harder (${\rm REG}(\nu_{i})\to\infty$) must approach the set of ill-conditioned distributions (${\rm RSE}(\nu_{i})\to 0$). In all examples we consider, this holds at least on bounded sets $\mathcal{Q}^{\prime}\subset\mathcal{P}_{2}$. The proof of the following elementary proposition appears in Appendix B.1.